OpenStudy (anonymous):

Think about a kayaking race, a round trip airplane flight, a shopper on an escalator, or a swimming competition. Choose an activity where a rational expression can represent the total or difference of time. For example, what is the time to go round trip on a hot air balloon ride? Use full sentences to describe the real world situation. Write the addition or subtraction rational expression to represent your situation. Simplify the expression, showing all the work or explaining the steps using complete sentences.

OpenStudy (anonymous):

please help, i don't undeerstand this

OpenStudy (ash2326):

Suppose you are in a swimming competition and the competition is for 5 laps. How is that?

OpenStudy (anonymous):

i dont understand how to make the problem

OpenStudy (ash2326):

OK, I'll explain you:)

OpenStudy (anonymous):

okay thank you:)

OpenStudy (ash2326):

just 2 minutes

OpenStudy (anonymous):

alrighty. :)

OpenStudy (ash2326):

Thanks for waiting

OpenStudy (ash2326):

Suppose I'm in a swimming competition. The competition consists of 5 laps. The competition began sometime back, and I've completed 2 laps. So how much of the race I've completed.

OpenStudy (anonymous):

no problem, your helping me.

OpenStudy (anonymous):

how would you put that into a math problem though?

OpenStudy (ash2326):

Out of 5 laps, I've completed 2 laps so I've completed \(\large \frac{2}{5}\)th of the race.

OpenStudy (ash2326):

Now I'm in the middle of the competition and I need to know how much more race is remaining. The full competition consists of 5 laps, so \[\frac{5}{5}=1\] I've completed \(\large\frac25\)th of the race, so \[\text{Competition remaining}=1-\frac25=\frac35\] SO \(\huge \frac35\)th race is remaining

OpenStudy (ash2326):

Did you understand how we saw a real life situation from a Mathematical view?

OpenStudy (anonymous):

I understand that, but in the lessom the problems are like this... Ex. \[4\over x-7 + 1\over x-3\]

OpenStudy (ash2326):

Let me think, how we could get like that

OpenStudy (anonymous):

Okay. Thank you.

OpenStudy (ash2326):

I'm here, thinking:)

OpenStudy (anonymous):

Okay:)

OpenStudy (ash2326):

Supposedly I've completed x laps!!! and I know 1/5 th race is remaining, so With x laps, fraction of race completed \[\frac x5\] \[\text{Full Race corresponds to }1\] so \[1-\frac x5= \frac15\] This can be used to evaluate the no. of laps completed in the competition:D